### 17 Mar 2026 v2

# On the separation Łojasiewicz exponents of real analytic sets in the real plane

Abstract.
The main aim of the paper is to give a formula for computing the separation Łojasiewicz exponents for two real analytic set germs via the Newton–Puiseux expansions of their defining functions. Moreover, we present an effective exponent for the case of two real algebraic sets in terms of their degrees.
Key words and phrases:
Łojasiewicz exponents, separation Łojasiewicz inequality, Newton polygon, Newton–Puiseux roots, effective Łojasiewicz inequality
1991 Mathematics Subject Classification:
14H20, 14B05, 32B05, 58K05
The first and second author’s research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.04-2024.09.
1. Introduction
The Łojasewicz inequalities impressively appeared in the 50s years of last century to solve the question of L. Schwartz for distribution division [13, 20, 21]. They have many relations and applications to other branches of mathematics, such as research on local singularities of analytic functions [26], the proof of Thom’s Gradient Conjecture [19], study of infinitely dimensional version in partial differential equations [3], applications in polynomial and tame optimization [11, 12], …
Let $A,B$ be real semi-analytic subsets of an open subset $U\subseteq\mathbb{R}^{n}$ with $0\in A\cap B$. Then we have the following inequality [21]:
  • There exist $C,r>0$ and $\beta\geq 1$ such that
    $d(x,A)+d(x,B)\geq Cd(x,A\cap B)^{\beta},\text{for all}\ \|x\|\leq r,$ (1)
    where $d(x,X)$ is the Euclidean distance of $x\in\mathbb{R}^{n}$ to the set $X$ ($d(x,X)=1$ if $X=\emptyset$).
  • The infimum of such exponents $\beta$ is called separation Łojasiewicz exponent of semi-analytic subsets $A$ and $B$ (at the origin), it is denoted by $\mathscr{L}(A,B)$.
Note that the separation Łojasiewicz exponent $\mathscr{L}(A,B)$ is a rational number [1, Corollary 2]. The Łojasiewicz exponent represents and compares the growths of functions on two sides of the inequalities [21].
The Łojasiewicz exponents are used to study some important problems, such as computing some topological invariants [23, 25, 26], investigating the singularity at infinity [8], the study of the separation of real algebraic sets [4, 18] proof of effective Nullstellensatz [14, 15], etc. There are many works providing computations and estimations of Łojasiewicz exponents [7, 9, 10, 5, 18, 23, 24].
In this paper, we study the separation Łojasiewicz exponents of the zeros of the real analytic functions in two variables. More precisely, assuming $f$ and $g$ are real analytic functions in two variables, we give a formula to compute the separation Łojasiewicz exponent $\mathscr{L}(f^{-1}(0),g^{-1}(0))$ in (1) in terms of the approximations of the Newton–Puiseux roots of $f\cdot g$ (see Definition 2.6 and Theorem 3.1). If $f$ and $g$ are polynomials, then we give an effective formula for the separation Łojasiewicz exponent.
2. Preliminaries
2.1. Newton polygon relative to an arc
In this section, we recall the technique of the Newton polygon relative to an arc or sliding-technique due to Kuo-Parusinski [17] (see also [10]). This technique plays a key role in this article. Let $f:(\mathbb{K}^{2},0)\to(\mathbb{K},0)$ ($\mathbb{K}$ is $\mathbb{C}$ or $\mathbb{R}$) be an analytic function germ. Suppose that $f$ is mini-regular in $x$ of order $m$, i.e. in the Taylor expansion of $f$:
$f(x,y)=f_{m}(x,y)+f_{m+1}(x,y)+\dots,$
we have $f_{m}(1,0)\neq 0$, where $f_{k}(x,y)$ is the homogeneous component of degree $k$.
For a Puiseux series
$x=\phi(y)=c_{1}y^{n_{1}/N}+c_{2}y^{n_{2}/N}+\cdots\in\mathbb{K}\{y^{1/N}\}$
with $N\leq n_{1}<n_{2}<\cdots$ being positive integers, where $c_{1}\neq 0$. We define the order of a Puiseux series $x=\phi(y)$:
$\mathrm{ord}\phi=\frac{n_{1}}{N}.$
The series $\phi$ is said to be real if all coefficients $c_{i}$ of $\phi$ are real. Let us define
$M(X,Y):=f(X+\phi(Y),Y):=\sum c_{ij}X^{i}Y^{j/N}.$
For each $c_{ij}\neq 0,$ let us plot a dot at $(i,j/N)$ in $\mathbb{R}^{2}$, and call it a Newton dot. The set of Newton dots is called the Newton diagram, They generate a convex hull, whose boundary is called the Newton polygon of $f$ relative to $\phi,$ to be denoted by $\mathbb{P}(f,\phi).$ Note that this is the Newton polygon of $M$ in the usual sense.
Algorithm 2.1 (The sliding).
  1. Input.
    An analytic function germ $f:(\mathbb{K}^{2},0)\to(\mathbb{K},0)$, mini-regular in $x$, and an initial Puiseux series $\phi_{0}(y)$. Set $k:=0$ and $\phi_{0}:=\phi$.
  2. Step 1.
    Define M_k(X,Y):=f(X+ϕ_k(Y),Y) =∑c^(k)_ijX^iY^j/N and consider the Newton polygon $\mathbb{P}(f,\phi_{k})$ of $M_{k}$. Let $(0,h_{k})$ be the lowest Newton dot on $X=0$ and $E_{k}$ be the compact edge of $\mathbb{P}(f,\phi_{k})$ having $(0,h_{k})$ as a vertex. Then ordf(ϕ_k(y),y)=h_k.
  3. Step 2.
    • If $\mathbb{P}(f,\phi_{k})$ has no Newton dot on $X=0$, then $f(\phi_{k}(y),y)=0$ and the algorithm stops with output
      $\phi_{\infty}(y):=\phi_{k}(y).$
    • If $\mathbb{P}(f,\phi_{k})$ has a Newton dot on $X=0$, choose a nonzero root $c_{k}$ of the edge polynomial
      $\mathcal{E}_{E_{k}}(z):=\sum_{(i,j/N)\in E_{k}}c^{(k)}_{ij}z^{i}$
      and set
      $\phi_{k+1}(y):=\phi_{k}(y)+c_{k}\,y^{\tan\theta_{E_{k}}},$
      where $\theta_{E_{k}}$ is the angle of the edge $E_{k}$.
  4. Step 3.
    Increase $k:=k+1$ and return to Step 1.
  5. Output.
    The sequence $\{\phi_{k}(y)\}_{k}$ has diagram $\phi_{0}\rightarrow\phi_{1}\rightarrow\cdots\rightarrow\phi_{k}\rightarrow\cdots$ which produces a Puiseux series $\phi_{\infty}$ satisfying
    $f(\phi_{\infty}(y),y)=0.$
The series $\phi_{\infty}$ is called a final result of the sliding of $\phi$ along $f$. The above algorithm is based on the following lemma.
Lemma 2.1.
Suppose that $\phi$ is not a root of $f=0$. Consider a series
$\psi:x=\phi(y)+cy^{\rho}+o(\rho),$
where $c\in\mathbb{K},0<\rho\in\mathbb{Q}$ and $o(\rho)$ denotes a Puiseux series of order greater than $\rho$. Then the following statements hold:
  • (i)
    If $\tan\theta_{E_{1}}<\rho$ or $\tan\theta_{E_{1}}=\rho$ and $\mathcal{E}_{E_{1}}(c)\neq 0$ then $\mathbb{P}(f,\phi)=\mathbb{P}(f,\psi),$ and, therefore $\mathrm{ord}f(\phi(y),y)=\mathrm{ord}f(\psi(y),y).$
  • (ii)
    If $\tan\theta_{E_{1}}=\rho$ and $\mathcal{E}_{E_{1}}(c)=0$ then $\mathrm{ord}f(\phi(y),y)<\mathrm{ord}f(\psi(y),y)$.
Proof.
For a detailed proof, we refer to [9]. In fact, the special case where $\psi:x=\phi(y)+cy^{\tan\theta_{E_{1}}}$ was proved in [9, Lemma 2.1]. The lemma is then deduced by applying the special case (infinitely) many times. ∎
We define
[…]
where $V_{f}$ denotes the zero locus $f(x,y)=0$ of $f$.
Lemma 2.2.
Suppose that $\phi$ is not a root of $f=0$. Then
  • (i)
    $\mathrm{ord}d(\phi,V_{f})\leq\tan\theta_{E_{1}}.$
  • (ii)
    Let $\psi$ be a series of form
    $\psi:x=\phi(y)+cy^{\rho}+o(\rho),\ c\in\mathbb{R}.$
    Suppose that $\rho>\mathrm{ord}d(\psi,V_{f})$, or $\rho\geq\mathrm{ord}d(\psi,V_{f})$ and $c$ is generic. Then
    $\mathrm{ord}d(\psi,V_{f})=\mathrm{ord}d(\phi,V_{f}).$
Proof.
(i) Assume by contradiction that $\rho:=\mathrm{ord}d(\phi,V_{f})>\tan\theta_{E_{1}}$, then taking a real Newton-Puiseux root $\beta(y)$ such that $\mathrm{ord}d(\phi,V_{f})=\mathrm{ord}(\phi(y)-\beta(y))$ we obtain
$\beta(y)=\phi(y)+cy^{\rho}+o(\rho),c\neq 0.$
By Lemma 2.1, $\mathrm{ord}f(\phi(y),y)=\mathrm{ord}f(\beta(y),y)=+\infty$, which is a contradiction.
(ii) Let $\beta$ be any Newton-Puiseux root of $f$. Then
$\psi(y)-\beta(y)=\phi(y)-\beta(y)+cy^{\rho}+o(\rho).$
Since $\mathrm{ord}\left(\psi(y)-\beta(y)\right)\leq\mathrm{ord}d(\phi,V_{f})\leq\rho$, it follows that
$\mathrm{ord}\left(\psi(y)-\beta(y)\right)=\mathrm{ord}\left(\phi(y)-\beta(y)\right).$
Therefore
$\mathrm{ord}d(\psi,V_{f})=\mathrm{ord}d(\phi,V_{f}).$
The following lemma is similar to [10, Proposition 2.3].
Lemma 2.3.
Assume that $\mathrm{ord}\phi(y)\geq 1$. Then
$d((\phi(y),y),V_{f})\simeq|y|^{\mathrm{ord}d(\phi,V_{f})},\text{ as }y\to 0$
Proof.
(cf. [16, Lemma 2.16]). Let us consider the following two cases.
If $f=0$ does not have any real Newton–Puiseux root, then $V_{f}=(0,0)$. Taking $\epsilon>0$ small sufficiently such that $|y|<\epsilon$, we have
$d((\phi(y),y),V)=d((\phi(y),y),(0,0))=\|(\phi(y),y)\|=\sqrt{(\phi(y))^{2}+y^{2}}.$
Since the assumption $\mathrm{ord}\phi(y)\geq 1$, then
$d((\phi(y),y),V)\simeq|y|^{1}\ \text{for all}\ |y|<\epsilon.$ (2)
If $f=0$ has some real Newton–Puiseux roots, then there exists $\epsilon>0$ sufficiently small such that $f(x,y)$ can be represented as follows
$f(x,y)=u(x,y)\prod_{j=1}^{k}(x-\beta_{i}(y)),$
where $u(x,y)$ has no real roots and $|y|<\epsilon$. Hence
$d((\phi(y),y),V)=\inf_{j}\|(\phi(y),y)-(\beta_{j}(y),y)\|=\inf_{j}|\phi(y)-\beta_{j}(y)|.$
It follows that
$\inf_{j}|\phi(y)-\beta_{j}(y)|\simeq|y|^{\max_{j}\{\mathrm{ord}(\phi(y)-\beta_{j}(y))\}}.$
Therefore, $d((\phi(y),y),V)\simeq|y|^{\mathrm{ord}d(\phi,V)}$ for all $|y|<\epsilon$. Combining this with (2), we deduce the proof of the lemma. ∎
2.2. Approximations of Puiseux series
Let $x=\gamma(y)$ be a Puiseux series in the following form:
$\gamma(y)=a_{1}y^{n_{1}/N}+a_{2}y^{n_{2}/N}+\cdots+a_{s-1}y^{n_{s-1}/N}+c_{s}y^{n_{s}/N}+\cdots,$
where $a_{i}\in\mathbb{R}$ and $c_{s}$ is the first non-real coefficient, if there is one. Let us replace $c_{s}$ by a generic real number $g,$ and call
$\gamma^{\mathbb{R}}(y):=a_{1}y^{n_{1}/N}+a_{2}y^{n_{2}/N}+\cdots+a_{s-1}y^{n_{s-1}/N}+gy^{n_{s}/N},$ (3)
a real approximation of $\gamma$. In case $s=+\infty,$ let $\gamma^{\mathbb{R}}=\gamma.$
Let $f\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be a real analytic function germ, which is mini-regular in $x$ of order $m$. Suppose that $\phi$ is not a Newton–Puiseux root of $f=0$ and $\gamma$ is a final result of the sliding of $\phi$ along $f$. Then
$\gamma(y)=\phi(y)+cy^{\tan\theta_{E_{1}}}+\text{higher order terms},$
where $\mathcal{E}_{H_{M}}(c)=0$ and $M(X,Y)=f(X+\phi(Y),Y)$. If $c\notin\mathbb{R}$, then by the definition of the real approximation (3), the real approximation $\gamma^{\mathbb{R}}$ of the series $\gamma$ also has the form
$\gamma^{\mathbb{R}}(y)=\phi(y)+gy^{\tan\theta_{E_{1}}},$ (4)
where $g\in\mathbb{R}$ is generic. For $f\in\mathbb{K}\{x,y\}$ which is regular in $x$, let $\mathcal{V}_{\mathbb{R}}(f)$ be the set of all real approximations of all the Newton-Puiseux roots of $f$.
Definition 2.4 ([23]).
Let $\phi$ be a Puiseux series $\phi(y)=\sum a_{i}y^{\alpha_{i}}$. For each positive real number $\rho$, the $\rho$-approximation of $\phi(y)$ is defined by series
$\sum_{\alpha_{i}<\rho}a_{i}y^{\alpha_{i}}+cy^{\rho},$
where $c$ is a generic real number.
Remark 2.5.
Let $\phi(y)$ be a non-real Puiseux series. Then the real approximation $\phi^{\mathbb{R}}(y)$ of $\phi(y)$ is the $\rho$-approximation of $\phi$, where $\rho$ is the smallest exponent occurring in $\phi$ with non-real coefficient.
Definition 2.6.
Let $\phi_{1},\phi_{2}$ be any two distinct Puiseux series. The series $\phi_{1,2}$ is called the approximation of $\phi_{1},\phi_{2}$ if it is the $\rho$-approximation of $\phi_{1}$ (also the $\rho$-approximation of $\phi_{2}$), where $\rho:=\text{ord}(\phi_{1}-\phi_{2})$. Let $f\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be a real analytic function germ, which is mini-regular in $x$. Let $x=\phi_{i}(y),i=1,\ldots,s$ be the set of the Newton-Puiseux roots of $f$. We denote by $\mathcal{V}_{a}(f)$ the set of all approximations $\phi_{i,j}$ of $\phi_{i}$ and $\phi_{j}$ such that $\phi_{i,j}$ is a real Puiseux series. Note that $\mathcal{V}_{\mathbb{R}}(f)\subset\mathcal{V}_{a}(f)$ if $f\in\mathbb{R}\{x,y\}$.
Example 2.7.
Let $\phi_{1}(y)=y^{\frac{3}{2}}-2y^{\frac{5}{2}}+3iy^{\frac{7}{2}}+\dots$ and $\phi_{2}(y)=y^{\frac{3}{2}}-2y^{\frac{5}{2}}-3y^{\frac{7}{2}}+iy^{\frac{9}{2}}+\dots$. Then, the approximation of series $\phi_{1}$ and $\phi_{2}$ is $\phi_{1,2}(y)=y^{\frac{3}{2}}-2y^{\frac{5}{2}}+gy^{\frac{7}{2}}$, where $g\in\mathbb{R}$ and generic.
3. Separation Łojasewicz exponent of two real analytic subsets
Let $f,g\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be reduced real analytic functions germs in two variables, which are mini-regular in $x$ of order $m$. Put $V_{f}=f^{-1}(0),V_{g}=g^{-1}(0)$. Then, by the Łojasiewicz inequality (1), there exist $C,r>0$ and $\beta\geq 1$ such that
$d(x,V_{f})+d(x,V_{g})\geq Cd(x,V_{f}\cap V_{g})^{\beta}\text{ for all}\ \|x\|\leq r.$ (5)
Let $\varphi\colon[0,\varepsilon)\to\mathbb{R}^{2}$ be a semi-algebraic analytic curve such that $\varphi(0)=0$. Then there is a rational number $\ell(\varphi)$ such that
$d(\varphi(t),V_{f})+d(\varphi(t),V_{g})\sim\left(d(\varphi(t),V_{f}\cap V_{g})\right)^{\ell(\varphi)}\text{ as }t\to 0.$
Let $\mathscr{L}(V_{f},V_{g})$ denote the separation Łojasiewicz exponent of $V_{f}$ and $V_{g}$. Using the Curve Selection Lemma (see [22]), we can easily prove the following equality
$\mathscr{L}(V_{f},V_{g})$ $=$ $\sup_{\phi}\ell(\phi),$ (6)
where the supremum is taken over all semi-algebraic analytic curves passing through the origin, which do not contain in $V_{f}\cap V_{g}$.
Suppose that the curve $\varphi$ is parametrized by a Puiseux series $x=\phi(y)$. We can see that
$\ell(\varphi)=\ell(\phi):=\frac{\min\{\mathrm{ord}d(\phi,V_{f}),\mathrm{ord}d(\phi,V_{g})\}}{\mathrm{ord}d(\phi,V_{\gcd(f,g)})}.$ (7)
We define the number
$\mathscr{L}_{+}(V_{f},V_{g}):=\max\{\ell(\gamma)|\gamma\in\mathcal{V}_{a}(f\cdot g)\},$ (8)
where the set $\mathcal{V}_{a}(f\cdot g)$ is defined in Definition 2.6. Let us denote by $\overline{f},\overline{g}$ the real analytic function germs defined by $\overline{f}(x,y):=f(x,-y)$ and $\overline{g}(x,y):=g(x,-y)$, and set $\mathscr{L}_{-}(V_{f},V_{g}):=\mathscr{L}_{+}(V_{\overline{f}},V_{\overline{g}})$. Now, we establish the formula for the separation Łojasiewicz exponent $\mathscr{L}(V_{f},V_{g})$.
Theorem 3.1.
Let $f,g\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be non-zero reduced real analytic function germs, which are mini-regular in $x$ of order $m$. Then
$\mathscr{L}(V_{f},V_{g})=\max\{\mathscr{L}_{+}(V_{f},V_{g}),\mathscr{L}_{-}(V_{f},V_{g})\},$
where $\mathscr{L}_{-}(V_{f},V_{g}):=\mathscr{L}_{+}(V_{\overline{f}},V_{\overline{g}})$ and $\mathscr{L}_{+}(V_{\overline{f}},V_{\overline{g}})$ is defined by (8).
First, we need some lemmas. Put $h:=\gcd(f,g)$, where $\gcd(f,g)$ is the greatest common division of $f$ and $g$. Let us consider the Newton polygons relative to any arc $x=\phi(y)$, they are $\mathbb{P}(f,\phi)$, $\mathbb{P}(g,\phi)$ and $\mathbb{P}(h,\phi)$ with the highest Newton edges $F_{1}$, $G_{1}$ and $H_{1}$, respectively. Assume that $\mathcal{E}_{F_{1}}$, $\mathcal{E}_{G_{1}}$ and $\mathcal{E}_{H_{1}}$ are their associated polynomials. Let $\theta_{F_{1}}$, $\theta_{G_{1}}$ and $\theta_{H_{1}}$ be the Newton angles of $F_{1}$, $G_{1}$ and $H_{1}$, respectively. It is easy to see that
$\min\{\tan\theta_{F_{1}},\tan\theta_{G_{1}}\}\geq\tan\theta_{H_{1}}.$ (9)
Lemma 3.2.
Suppose that $\ell(\phi)>\mathscr{L}_{+}(V_{f},V_{g})$. Then
  • (i)
    $\tan\theta_{F_{1}}=\tan\theta_{G_{1}}.$
  • (ii)
    The polynomial $\mathcal{E}_{F_{1}}\mathcal{E}_{G_{1}}$ has only one root, which is a real value.
Proof.
By contradiction we assume that $\tan\theta_{F_{1}}\neq\tan\theta_{G_{1}}.$ Without loss of generality suppose that $\tan\theta_{F_{1}}>\tan\theta_{G_{1}}.$ Let $\phi_{1,\infty}$ and $\phi_{2,\infty}$ be final results of sliding $\phi$ along $f$ and $g$, respectively. Then we have
$\phi_{1,\infty}(y)$ $=\phi(y)+\sum_{i\geq 1}a_{i}y^{\alpha_{i}},a_{1}\neq 0$ (10)
$\phi_{2,\infty}(y)$ $=\phi(y)+\sum_{i\geq 1}b_{i}y^{\beta_{i}},b_{1}\neq 0,$ (11)
where $\tan\theta_{F_{1}}=\alpha_{1}<\alpha_{2}<\ldots$ and $\tan\theta_{G_{1}}=\beta_{1}<\beta_{2}<\ldots$.
Let $\phi_{1,2}$ be the approximation of $\phi_{1,\infty}$ and $\phi_{2,\infty}$. Then $\phi_{1,2}\in\mathcal{V}_{a}(fg)$ and
$\phi_{1,2}(y)=\phi(y)+cy^{\tan\theta_{G_{1}}}+o(\tan\theta_{G_{1}}),$
where $c\in\mathbb{R}$ is generic. It follows from Lemma 2.2 that
$\mathrm{ord}d(\phi_{1,2},V_{g})=\mathrm{ord}d(\phi,V_{g})\text{ and }\mathrm{ord}d(\phi_{1,2},V_{h})=\mathrm{ord}d(\phi,V_{h}).$ (12)
We will show that $\ell(\phi_{1,2})=\ell(\phi)$ by considering the following two cases.
  • Case 1: $\mathrm{ord}d(\phi,V_{f})\leq\tan\theta_{G_{1}}$. Then $\mathrm{ord}\ d(\phi_{1,2},V_{f})=\mathrm{ord}\ d(\phi,V_{f})$ due to Lemma 2.2. Hence $\ell(\phi_{1,2})=\ell(\phi)$.
  • Case 2: $\mathrm{ord}d(\phi,V_{f})>\tan\theta_{G_{1}}$. Take a real Newton-Puiseux root $\gamma$ of $f$ satisfying
    $\mathrm{ord}(\phi-\gamma)=\mathrm{ord}\ d(\phi,V_{f}).$
    Since
    $\phi_{1,2}(y)-\gamma(y)=\phi(y)-\gamma(y)+cy^{\tan\theta_{G_{1}}}+o(\tan\theta_{G_{1}}),$
    it follows that $\mathrm{ord}d(\phi_{1,2}-\gamma)=\tan\theta_{G_{1}}$. Therefore,
    $\mathrm{ord}d(\phi_{1,2},V_{f})\geq\tan\theta_{G_{1}}\geq\mathrm{ord}d(\phi_{1,2},V_{g}).$ (13)
    Hence,
    $\ell(\phi_{1,2})$ $=\frac{\min\{\mathrm{ord}d(\phi_{1,2},V_{f}),\mathrm{ord}d(\phi_{1,2},V_{g})\}}{\mathrm{ord}d(\phi_{1,2},V_{h})}$
    $=\frac{\mathrm{ord}d(\phi,V_{g})}{\mathrm{ord}d(\phi,V_{h})}=\ell(\phi).$
The equality $\ell(\phi_{1,2})=\ell(\phi)$ contradicts the assumption that $\ell(\phi)>\mathscr{L}_{+}(V_{f},V_{g})$. This proves (i).
(ii) By (i) one has $\rho=\tan\theta_{F_{1}}=\tan\theta_{G_{1}}$. Then $\mathcal{E}_{F_{1}}\mathcal{E}_{G_{1}}$ is the polynomial associated with the highest Newton edge of $\mathbb{P}(fg,\phi)$ with the angle $\rho$. Suppose by contradiction that $\mathcal{E}_{F_{1}}\mathcal{E}_{G_{1}}$ has two distinct roots $c_{1},c_{2}$. Let $\varphi_{i}=\phi+c_{i}y^{\rho}$ for each $i=1,2$, and let $\varphi_{i,\infty}$ be a final result of sliding $\varphi_{i}$ along $fg$. Let $\varphi_{1,2}$ be the approximation of $\varphi_{1,\infty}$ and $\varphi_{2,\infty}$. Then $\varphi_{1,2}\in\mathcal{V}_{a}(fg)$ and
$\varphi_{1,2}(y)=\phi(y)+cy^{\rho}+o(\rho),$
where $c\in\mathbb{R}$ is a generic number. Note that
$\tan\theta_{H_{1}}\leq\tan\theta_{F_{1}}=\tan\theta_{G_{1}}=\rho.$
It follows from Lemma 2.2 that
$\mathrm{ord}d(\varphi_{1,2},V_{f})=\mathrm{ord}d(\phi,V_{f}),\ \mathrm{ord}d(\varphi_{1,2},V_{g})=\mathrm{ord}d(\phi,V_{g})$
and
$d(\varphi_{1,2},V_{h})=\mathrm{ord}d(\phi,V_{h}).$
Hence $\ell(\phi_{1,2})=\ell(\phi)$, which is a contradiction. The lemma is proved. ∎
Proof of Theorem 3.1.
By the formula (6), it is obvious that
$\mathscr{L}(V_{f},V_{g})\geq\max\{\mathscr{L}_{+}(V_{f},V_{g}),\mathscr{L}_{-}(V_{f},V_{g})\}.$
Suppose for contradiction that
$\mathscr{L}(V_{f},V_{g})>\max\{\mathscr{L}_{+}(V_{f},V_{g}),\mathscr{L}_{-}(V_{f},V_{g})\}.$
Then there is a real analytic arc $\phi$ passing through the origin and not lying the $x$-axis such that
$\ell(\phi)>\max\{\mathscr{L}_{+}(V_{f},V_{g}),\mathscr{L}_{-}(V_{f},V_{g})\}$ (14)
and the parameter forms of $\phi$ is
$(x=\phi(t),y=t)\ \text{or}\ (x=\phi(t),y=-t),\ \text{where}\ \phi(t)\in\mathbb{R}\{t^{1/N}\}.$
Without loss of generality, assume that $\phi$ can be parametrized by $(x=\phi(t),y=t)$. By Lemma 3.2, then $\theta_{F_{1}}=\theta_{G_{1}}$ and the polynomial $\mathcal{E}_{F_{1}}\mathcal{E}_{G_{1}}$ has a unique root $c_{0}\in\mathbb{R}$. Define
$\tilde{\phi}(y):=\phi(y)+c_{0}y^{\tan\theta_{F_{1}}}.$
Claim 3.1.
We have
$\ell(\tilde{\phi})\geq\ell(\phi).$
Proof.
We first see that
$\min\{\mathrm{ord}d(\phi,V_{f}),\ \mathrm{ord}d(\phi,V_{g})\}>\mathrm{ord}d(\phi,V_{f}\cap V_{g})=\mathrm{ord}d(\phi,V_{h}),$
because $\ell(\phi)>1$. Then
$\tan\theta_{F_{1}}\geq\mathrm{ord}d(\phi,V_{f})>\mathrm{ord}d(\phi,V_{h}).$
Applying Lemma 2.2 and its proof, one has
$\mathrm{ord}d(\tilde{\phi},V_{h})=\mathrm{ord}d(\phi,V_{h}),\ \mathrm{ord}d(\tilde{\phi},V_{f})\geq\mathrm{ord}d(\phi,V_{f})\text{ and }\mathrm{ord}d(\tilde{\phi},V_{g})\geq\mathrm{ord}d(\phi,V_{g}).$
Hence $\ell(\tilde{\phi})\geq\ell(\phi).$
[…]
Claim 3.2.
If $\tilde{\phi}$ is not a Newton–Puiseux root of $h=0$, then
  1. (1)
    $\tan\theta_{\widetilde{F}_{1}}=\tan\theta_{\widetilde{G}_{1}}$;
  2. (2)
    The polynomial $\mathcal{E}_{\widetilde{F}_{1}}\mathcal{E}_{\widetilde{G}_{1}}$ has only real root.
Let $\phi_{0}:=\phi$ and
$\phi_{1}(y)=\tilde{\phi}(y)=\phi_{0}(y)+c_{0}y^{\tan\theta_{F_{1}}}.$
Then $\phi_{1}$ is the unique sliding of $\phi_{0}$ along $f$ (as well as $g$ and $h$). Applying this process, we obtain a sequence
$\phi\to\gamma_{1}\to\gamma_{2}\to\dots\to\gamma_{n}\to\dots\to\phi_{\infty},$ (15)
where $\phi_{k+1}$ be the unique sliding of $\phi_{k}$ along $h$ for each $k\geq 1$. The series $\phi_{\infty}$ is a real Newton-Puiseux root of $h$. It yields the following identities
$\mathrm{ord}d(\phi,V_{f})=\mathrm{ord}d(\phi,V_{g})=\mathrm{ord}d(\phi,V_{h})=\mathrm{ord}(\phi-\phi_{\infty})=\tan\theta_{F_{1}}.$
This implies $\ell(\phi)=1$, which is a contradiction. So, the theorem is proved. ∎
4. Effective estimation for separation Łojasiewicz exponent
Suppose that $f$ and $g$ are polynomials of two real variables, then there is the following effective version of the separation Łojasiewicz exponent.
Theorem 4.1.
Let $f$ and $g$ be polynomials of two real variables such that their degrees are not greater than $d$ and $f(0)=g(0)=0$. Then there exist positive numbers $C$ and $r$ such that
$d(x,V_{f})+d(x,V_{g})\geq Cd(x,V_{f}\cap V_{g})^{\frac{(2d-1)^{2}+1}{2}},\ \text{for all}\ \|x\|\leq r.$ (16)
Proof.
(cf. [18]) Put $F:=f^{2}+g^{2}$, then $\deg F\leq 2d$, where $d=\max\{\deg f,\deg g\}$. Moreover, $V_{F}=V_{f}\cap V_{g}$.
If $V_{f}\subset V_{g}$, then $\mathscr{L}(V_{f},V_{g})=1$. Assume that $V_{f}\not\subset V_{g}$. Applying the classical Łojasiewicz inequality to $F$, there exists $\alpha,c>0$ such that
$|F(x)|\geq cd(x,V_{F})^{\alpha},\forall x\in U\cap(V_{f}\setminus V_{g}),$
where $U\subset\mathbb{R}^{2}$ is a neighbourhood of $(0,0)$. Let $\widetilde{\mathcal{L}}(F)$ be the classical Łojasiewicz exponent of the polynomial $F$, then, by [23, Theorem 4.4], $\widetilde{\mathcal{L}}(F)\leq(2d-1)^{2}+1$. Hence, the effective optimal version of classical Łojasiewicz inequality is
$|F(x)|\geq cd(x,V_{F})^{(2d-1)^{2}+1},\forall x\in U\cap(V_{f}\setminus V_{g}),$
Note that $x\in U\cap(V_{f}\setminus V_{g})$, then
$|F(x)|=|g(x)|^{2}.$ (17)
Therefore,
$|g(x)|\geq cd(x,V_{f}\cap V_{g})^{\frac{(2d-1)^{2}+1}{2}},\forall x\in U\cap(V_{f}\setminus V_{g}),$ (18)
It is easy to see that there exists $C,C^{\prime}>0$ such that
$Cd(x,V_{g})\geq C^{\prime}|g(x)|.$ (19)
Combining (17) with (19), we obtain
$d(x,V_{g})\geq Cd(x,V_{f}\cap V_{g})^{\frac{(2d-1)^{2}+1}{2}},\forall x\in U\cap(V_{f}\setminus V_{g}).$
This implies inequality (16). ∎
Remark 4.2.
The exponent in inequality (16) is sharper than the exponent in [18, Corollary 8] in the case of two variables, where $0\in\mathbb{R}^{2}$ is not necessarily isolated point of $V_{f}\cap V_{g}$.